OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (7, 4); Slope: 6 y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7)

OpenStudy (ccswims):

point-slope form y - y1 = m(x - x1)

OpenStudy (anonymous):

ok i got this

OpenStudy (anonymous):

y – 4 = 6(x – 7)

OpenStudy (anonymous):

the answer is y - 4 = 6(x-7)

OpenStudy (anonymous):

there you go all set!!!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

@knewitxx your welcome